# Flying Bird Object Detection Algorithm in Surveillance Video

### v1

@@ V Conclusion
In this paper, we have conducted an in-depth study of flying bird objects in surveillance videos. It is found that the characteristics of the flying bird object in the surveillance video are not obvious in a single-frame image, and the flying bird object is small and asymmetric in most cases. According to these characteristics, a method, FBOD-SV, for detecting flying birds in surveillance video is proposed. Firstly, aiming at the problem that the features of the single frame of the flying bird object in the surveillance video are not obvious, a new feature aggregation module, the Co-Attention-FA module, is designed to aggregate the features of the flying bird object according to the correlation of the flying bird object on multiple consecutive frames of images. Secondly, in view of the small size of flying birds in surveillance videos, in most cases, the FBOD-Net is designed, which is first down-sampling and then up-sampling. A large feature layer that combines fine spatial information and large receptive field information is used to detect special multi-scale flying birds. Then, according to the asymmetric and regular characteristics of the flying birds in the surveillance video, the SimOTA dynamic label allocation method is simplified, and the SimOTA dynamic label allocation strategy for One-Category object detection, SimOTA-OC dynamic label allocation strategy, is proposed to solve the difficult problem of label allocation caused by the irregular flying birds. Finally, a series of quantitative and qualitative experiments are designed to prove the effectiveness of the proposed bird object detection method in surveillance video, and the following conclusions are drawn:

### v3

@@ V Conclusion
In this paper, we have conducted an in-depth study of flying bird objects in surveillance videos. It is found that the characteristics of the flying bird object in the surveillance video are not obvious in a single-frame image, and the flying bird object is small and asymmetric in most cases. According to these characteristics, a method, FBOD-SV, for detecting flying birds in surveillance video was proposed. Firstly, addressing the issue of non-obvious features in single frames of flying bird objects in surveillance video, a new feature aggregation module, namely the Co-Attention-FA module, was devised. This module aggregates the features of the flying bird object by exploiting the correlation across multiple consecutive frames of images. Secondly, considering the multi-scale (particularly small size in most cases) character of flying birds in surveillance videos, the FBOD-Net was designed with a structure that first down-samples and then up-samples. A large feature layer that combines fine spatial information and large receptive field information is used to detect the flying birds. Then, considering the asymmetric and irregular characteristics of flying birds in surveillance video, we simplified the SimOTA dynamic label allocation method and proposed the SimOTA-OC dynamic label allocation strategy for One-Category object detection. This strategy aims to address the issue of label allocation posed by the irregular flying patterns of birds. Finally, a series of quantitative and qualitative experiments were designed to prove the effectiveness of the proposed bird object detection method in surveillance video, and the following conclusions are drawn: